OpenStudy (love_to_love_you):

Write the equation in slope-intercept form. What are the slope and y-intercept? -5x-12y=11

jigglypuff314 (jigglypuff314):

slope intercept form has the y by itself on one side so try to get y all alone first

OpenStudy (love_to_love_you):

I think it is d

OpenStudy (love_to_love_you):

a) y = -5/12x+11/12; slope: -5/12; y-int.: 11/12 b) y = 5/12x-11/12; slope: 5/12; y-int.: -11/12 c) y = -5/12x-11/12; slope: -5/12; y-int.: -11/12 d) y = 5/12x+11/12; slope: 5/12; y-int.: 11/12

jigglypuff314 (jigglypuff314):

mmm be careful ^_^ -5x - 12y = 11 +5x +5x -12y = 5x + 11 /-12 /-12 /-12 y = -5/12 x - 11/12

OpenStudy (love_to_love_you):

oh. so is it c?

jigglypuff314 (jigglypuff314):

yes :) but do you see where you got it wrong? :P

OpenStudy (love_to_love_you):

I accidentally thought that when you divided two negatives it would be positive.

jigglypuff314 (jigglypuff314):

well yeah, but you weren't dividing two negatives here

OpenStudy (love_to_love_you):

I know that's why I got it wrong hahaha sorry

jigglypuff314 (jigglypuff314):

it's alright ^_^